OpenStudy (anonymous):

Using the graph, which of these ordered pairs is the best estimate of the solution of the system y = 3x + 2 and y = -1/4 x + 1?

OpenStudy (anonymous):

A. (0.5, 0.5) B. (-0.3, 1.1) C. (1.2, 0.8) D. (0.8, -0.5)

OpenStudy (dumbcow):

set the 2 equations equal and solve for x 3x+2 = -1/4x +1

OpenStudy (anonymous):

So basically just put the numbers in for x, and whichever equation equals the same for both, is my answer?

OpenStudy (anonymous):

???

OpenStudy (dumbcow):

yeah you could do that too :)

OpenStudy (anonymous):

Okay thanks! :)

OpenStudy (anonymous):

I tried plugging in all the numbers for x and none of them equal the same.. or am I doing the wrong thing for this problem? I don't get it.

OpenStudy (dumbcow):

Try plugging in x = -0.3 again 3(-.3) +2 = ? -1/4(-.3) +1 = ?

OpenStudy (dumbcow):

there is some rounding error

OpenStudy (anonymous):

3(-.3)+2 = 1.7 -1/4(-.3)+1 = 0.7

OpenStudy (anonymous):

So now what?

OpenStudy (dumbcow):

3*(-.3) = -.9

OpenStudy (anonymous):

Okay but my answer choices available are: A. (0.5, 0.5) B. (-0.3, 1.1) C. (1.2, 0.8) D. (0.8, -0.5)

OpenStudy (hoblos):

you must solve it graphically so you only need to find the point of intersection between the two lines y = 3x + 2 & y = -1/4 x + 1 from the graph then the solution is (-0.3, 1.1)

OpenStudy (anonymous):

Oh!!!!! Well that would've helped me a long time ago. Lol thank you so much hoblos! :) It means a lot! :)

OpenStudy (hoblos):

my pleasure XD